OpenStudy (anonymous):

Help with: cos^2x - (sin^2x)(cos^2x)

OpenStudy (anonymous):

Maybe divide the cos^2x's out?

hartnn (hartnn):

you mean factor out cos^2 x ... yes.

OpenStudy (anonymous):

Which would leave you with 1 - sin^2x, correct? Except that's not one of the answer choices.

hartnn (hartnn):

then use the identity \(\sin^2 x + \cos^2 x = 1\)

hartnn (hartnn):

so whats 1- sin^2 x = ...? from the above identity?

OpenStudy (anonymous):

cos^2x, but that's not an answer choice either, lol.

hartnn (hartnn):

there is a factored out cos^2 x too, right? :)

hartnn (hartnn):

\(\Large \cos^2 x \times \cos^2 x = ...?\)

OpenStudy (anonymous):

cos^4x, that is an actual choice. Thank you!

hartnn (hartnn):

i will write out the steps, \(\Large \cos^2 x - \sin^2 x \cos^2 x = \cos^2 x (1- \sin^2 x ) \\ \large = \cos^2 x \times \cos^2 x = \cos^4 x \) and welcome ^_^